OpenStudy (anonymous):

Help I have an assigment due at midnight and I am lost. I can get the answers in decimal but not sure how to get in fraction. Evaluate the following. Give your answer as fractions; NO DECIMALS. If sin t = 3/5 and the terminal point of is in quadrant II, then fill in the following values. cos t = tan t = cot t= sec t = csc t =

OpenStudy (anonymous):



OpenStudy (anonymous):

you have a 3-4-5 right triangle, so take the ratios, and make sure to know that in quadrant II sine is positive, cosine is negative, tangent is negative, etc

OpenStudy (anonymous):

\[\cos(t)=-\frac{4}{5}\] \[\tan(t)=-\frac{3}{4}\] etc

OpenStudy (anonymous):

how did your know it was a 3-4-5 triangle??

OpenStudy (anonymous):

sin=opposite/hypotenuse so he knows the longest side is 5 and one of the sides is 3 and using pythagoreans theorem he deduced the last side is 4